# Physics-Enhanced Deep Surrogate for the Phonon Boltzmann Transport Equation

### v1

@@ 1 Introduction
Controlling heat flow at the nano-scale is essential for microelectronics, thermoelectrics, and energy-conversion technologies [Cahill2003, Vineis2010, Tang2010, wang2016all]. At these length scales the established modeling framework is the phonon Boltzmann Transport Equation (BTE), which resolves the phonon distribution in real and momentum space but is far more expensive to solve than classical diffusive models [Peierls1929, ziman2001electrons, Cahill2003, Chen2005, cahill2014nanoscale]. This computational cost is especially prohibitive in inverse design settings, where many forward solves over complex porous geometries are required for optimization [Evgrafov2009, Evgrafov2011]. Inverse design for nano-scale heat transport was first introduced in [romano2022inverse]. This approach performed density-based topology optimization [gersborg2006topology, lundgaard2018density] using differentiable phonon transport simulations and a novel interpolation technique to map the density of the material into an effective transmission coefficient. Despite this advance, a computational bottleneck arises in scenarios where multiple optimization runs need to create multiple target conductivities, because the topology optimization needs to be rerun from scratch for each target, which can become prohibitive due to the cost of solving the BTE.
@@
To address the computational costs, data-driven surrogates aim to accelerate the simulation and optimization of typically high-dimensional, design parameters search [GarciaEsteban2021, karniadakis2021physics]. In particular, machine learning surrogates hold the promise to learn the mapping from design parameters (i.e. material structure) to a desired low-dimensional target property (i.e. effective thermal conductivity) and dramatically reduce computational expense for the evaluation of the property. In addition, they offer a continuous relaxation of the parameters and the gradient may be computed efficiently via automatic differentiation [griewank2008evaluating, baydin2018automatic]. However, purely data-driven models come with the cost of the training data and are limited by the curse of dimensionality that require large training sets that grow exponentially with the number of variables [bishop2006pattern, hastie2009elements], causing an upfront, potentially large, training cost. Moreover, surrogate models may generalize poorly outside the generated data distribution, being often unreliable at extrapolation compared to interpolation [raissi2019physics]. Hybrid models from scientific machine learning (SciML) aim at reducing this data need by incorporating domain knowledge inside the model. Scientific and physics-informed approaches (SciML, PIML) have been used to create faster approximations [karniadakis2021physicsinformed, peurifoy2018nanophotonic, So2020DeepID, Kiarashinejad2020DimRed, kiarashinejad2020knowledge, kiarashinejad2019deep, An2019Objective, An2020Deep]. These include physics-informed networks [raissi2019physics, lu2021physics], solver corrections [mcbane2021component, chung2024scaled, zanardi2024scalable, choi2025defining, dresdner2022learning], and operator learning [li2020fourier, kovachki2023neural, lu2021learning]. Many of these models follow physical rules by including symmetries [cohen2016group], structural patterns [baddoo2023physics], or geometric information [bronstein2017geometric, gao2022physics], or by directly embedding solvers in the learning process  [Pestourie_Mroueh_Rackauckas_Das_Johnson]. This approach has enabled major results in protein folding [jumper2021highly] and weather forecasting [pathak2022fourcastnet, kochkov2024neural]. Physics-Informed Neural Networks (PINNs)  [raissi2019physics] have been extensively applied to diffusive  [Mishra2020] and ballistic heat transfer [li2021physicsinformedneuralnetworkssolving, li2022physics, li2023physics, lin2025monte] and to inverse problems in the diffusive  [Cai2021, Billah2023, Qian2023] and ballistic regime  [zhou2025physics]. Focusing on nano-scale transfer, in  [li2021physicsinformedneuralnetworkssolving, li2022physics, zhou2023physics, lin2025monte], PINNs are trained minimizing the sum of the residual with respect to the mode-resolved BTE and its boundary conditions at different points accross the domain (data-free ML). Recently, Ref. [zhou2025physics] proposed an architecture that extends to identifying some of the PDE parameters in an inverse problem setting. However, PINNs are solvers that learn a PDE solution rather than surrogate models that learn the parameterized function of a property. PINNs are typically not parameterized, but previous work introduced one physical parameter to the solver [li2021physicsinformedneuralnetworkssolving, li2022physics, zhou2023physics, lin2025monte] ––– the characteristic length which is hard to define and compute for complex geometries. In contrast, our surrogate model is geared towards design with twenty-five geometry parameters. Multi-fidelity (MF) DeepONets [lu2021learning] learn a parameterized PDE solver from data, integrating low-fidelity simulations with sparse high-fidelity data [lu2022multifidelity]. This approach improved predictive accuracy and reduced the cost of data generation. However, it required 1000 high-fidelity data, and the low-fidelity model still required 10,000 approximate BTE solves to work, which remained computationally expensive. In contrast, we need only a few hundred high-fidelity data and use the much cheaper macroscopic approximation of Fourier equation inside our model. Note that operator learning techniques like DeepONet reduce to a single neural network (one of our baselines) when used as a parameterized surrogate model [Pestourie_Mroueh_Rackauckas_Das_Johnson].
@@
We developed a SciML surrogate for the steady-state Boltzmann Transport Equation based on Physics Enhanced Deep Surrogate (PEDS) [Pestourie_Mroueh_Rackauckas_Das_Johnson]. PEDS combines a low-fidelity solver, enforcing physical behavior, with a neural network that learns the low-fidelity solver input that makes it accurate for a target property. The cheap solver provides an inductive bias [Mitchell1980] with a relaxed version of the physics (Sec. 2). In our case, we employ a differentiable Fourier simulator as the low-fidelity approximation of the BTE. To our knowledge, this is the first successful implementation of a multifidelity approach that leverages the diffusion equation for BTE. Although when computing the thermal conductivity it has a fractional error of up to 600%, the Fourier solver is $\approx$ 2300 times faster to compute (11000 times faster considering its strong batch-parallelism) and its inductive bias accelerates training and improves the generalization in out-of-distribution regions, resulting in three to four orders of magnitude accelerations of PEDS compared to BTE. Solving one high-fidelity BTE takes around 3 minutes on 4 CPUs for our illustrative example, so training costs are dominated by data simulation costs, and our method aims at reducing the data need while retaining accuracy. Combined with active learning [pestourie2020active], PEDS needs only 300 data points to achieve a 5.00% target fractional error (Sec.  4), dominated by fabrication error and sufficient for PDE-constrained optimization purposes (Sec.  4.2). The proposed surrogate outperforms a purely data-driven baseline, improving the model test set fractional error of up to 76% relatively to the data-driven baseline for the same number of training points (Sec.  4). Our current inverse design pipeline enables fast PDE evaluations for thermal conductivity design tasks. We achieve an average design error of $\sim$4% on 8 example design objectives, dominated by the error of the fabrication process. Thanks to the data efficiency of PEDS and Active Learning (AL), we fully amortize the training costs with only four designs.
@@ 2 An Overview of Physics Enhanced Deep Surrogates
Figure 1: PEDS Diagram The main panel illustrates the PEDS workflow: starting from a vector of topology parameters $G$, a neural-network generator $\text{generator}_{NN}$ produces an updated, coarse topology that encodes a learned, non-linear transformation of the design space. This generated topology is then combined with a coarsified representation of the original geometry via a linear combination; the mixing coefficient $w_{\phi}\in[0,1]$ is itself learned as a function of the geometry parameterization so that the model exploits the original design and the network-generated topology in a data- and physics-informed way. The resulting modified geometry is passed to a computationally cheap, low-fidelity version of the physics. In our case this is the Fourier solver (depicted here with a cartoon image of Ludwig Boltzmann) which evaluates the approximate physics and outputs the target macroscopic property. The top right inset highlights how the training dataset is constructed: high-fidelity ground truth labels are produced by running expensive Boltzmann Transport Equation simulations using OpenBTE (represented here by a historical photograph of Ludwig Boltzmann, courtesy of the University of Frankfurt). The top left insets represent the pure-data driven baselines that we considered, a Gaussian Process with a composite kernel and an MLP.
@@
PEDS is a scientific machine learning model comprising a computationally cheap low-fidelity solver paired with a neural network generator [Pestourie_Mroueh_Rackauckas_Das_Johnson], as in Eq. 1 and illustrated in Fig. 1. The neural network applies a nonlinear transformation to the design-parameter space; the transformed parameters are passed to the low-fidelity solver, which approximates the governing physics and returns the target property. In this sense, PEDS belongs to the input-space representation machine learning field [koziel2008space, marzban2025inverse] and is similar to neural space mapping [bakr2000neural] or coarse/fine grid mapping [feng2019coarse], with the relevant differences that the output parametrization is not the same as the initial and that in more classic input-space machine learning there is no mixing of the generated input with an input from field knowledge. The low-fidelity solver may be the same numerical method as the high-fidelity PDE solver, run at lower spatial resolution or with higher tolerance, or it may incorporate deliberate simplifications of the physics (for example, by linearizing a nonlinear term). This low-fidelity solver can produce large errors in the target output (in our case over 200 % on average, with up to almost 700% fractional error, as shown in Figure 2), but it is orders of magnitude faster than the high-fidelity model while qualitatively preserving at least some of the underlying physics (e.g. the boundary conditions). To incorporate further physical bias, the geometry fed to the approximate solver can be a linear combination of the generated topology and the original geometry, possibly coarsified to match the generated topology dimension. The coefficient of the convex linear combination, from now on referred to as the mixing coefficient, can be learned as a function of the input geometry. The modified geometry can also enforce physically sound biases like symmetries. The neural network weights are learned contextually using backpropagation and the adjoint simulation of the PDE. PEDS has previously shown great improvements in terms of accuracy and data efficiency compared to other deep parametric models and classical surrogates in similar tasks, including surrogate modeling for diffusion and diffusion-reaction equations, and for the more complex Maxwell’s equations [Pestourie_Mroueh_Rackauckas_Das_Johnson]. PEDS is defined as
@@
$\kappa\approx f_{lowfid}\left(w_{\phi}\text{generator}_{NN}(G)+(1-w_{\phi})\text{downsample}(G)\right),$ (1)
@@
where $G$ parameterizes the surrogate model input geometry, $f_{lowfid}$ is the low-fidelity solver, $w_{\phi}$ the mixing coefficient, $\text{generator}_{NN}$ the neural generator of the solver input, and downsample the solver input from field knowledge. PEDS fits greatly the Boltzmann Transport Equation, as the mixing coefficient $w_{\phi}$ dapts predictions between macro- and nano-scale effective conductivities, naturally capturing the smooth transition from diffusive to ballistic transport. Moreover, there is wide literature converging to a clear choice for the solver to approximate the physics: the Fourier equation, also known as heat conduction equation or Poisson equation. The Fourier Equation is the macro-scale approximation of the BTE and it can be derived through simplification of the BTE. When the average conductivity approaches the bulk conductivity (i.e. assuming uniformity of the material and absence of pores), the BTE scalar temperature field approaches the Fourier field (see (c) of figure 2). The Fourier Equation meets the inclusion criteria defined in [Pestourie_Mroueh_Rackauckas_Das_Johnson] , since uniform geometries encompass the full range of effective conductivities described by the BTE. Nonetheless, Fourier by itself overestimates the effective thermal conductivity of the material, especially in the smaller $\kappa$ regime dominated by ballistic transport.
@@
The neural network parameters and mixing coefficient are learned jointly in an end-to-end manner, making a differentiable low-fidelity solver essential. To efficiently compute the gradients of the PDE solution with respect to multiple design parameters, we exploit the adjoint simulation, also known as reverse-mode automatic differentiation [baydin2018automatic] or backpropagation, depending on the reader’s background. Instead of using system inversion or forward numerical differentiation, gradients are obtained by solving the adjoint equation, an auxiliary PDE that captures efficiently the solution’s sensitivity to design parameters. Obtaining the gradients amounts only to solving another system similar to the forward system. For self-adjoint operators, like Fourier with periodic boundary conditions, the adjoint system is identical to the original, and further computational efficiencies can be achieved depending on the re-usability of the forward solution strategy. Training and inference were performed on Intel Core i5 quad-core processors (2.3 GHz, 2017 generation) using CPU execution only.
@@
Figure 2: Comparison between the Fourier solver and the BTE. On the left we show the qualitative difference between the temperature fields obtained with (a) the ballistic BTE and (b) the diffusive Fourier model for the same geometry and problem setup. The domain is a square of size 100 nm, containing embedded square pores of size 10 nm, with periodic boundary conditions applied along both the x and y directions. To impose a temperature gradient, a bias of +1 K is enforced between the top and bottom boundaries along the y axis. In the BTE solution, the temperature field exhibits clear non-diffusive effects, with strongly distorted isothermal contours around the pores that highlight the influence of boundary scattering and ballistic transport. In contrast, the Fourier model produces a smooth field with symmetric and evenly spaced isothermal contours. The effective thermal conductivity $\kappa$ obtained by the Fourier temperature field is 105 W/mK, while it is 23 W/mK for the BTE field. On the right we present a quantitative comparison of the computed effective thermal conductivity for 500 nanoporous designs in our test set (c). The Fourier model systematically overestimates the thermal conductivity when compared with the BTE results (see the line x = y for reference). The overestimation is most severe for designs where the BTE conductivity is small, corresponding to cases dominated by ballistic transport, with fractional errors exceeding 700% in the most extreme examples and an average discrepancy of about 220%. As the effective conductivity increases, approaching the diffusive regime, the discrepancy between Fourier and BTE predictions decreases and the fractional error drops accordingly.

### v2

@@ 1 Introduction
Controlling heat flow at the nano-scale is essential for microelectronics, thermoelectrics, and energy-conversion technologies [11, 91, 89, 93]. At these length scales the established modeling framework is the phonon Boltzmann Transport Equation (BTE), which resolves the phonon distribution in real and momentum space but is far more expensive to solve than classical diffusive models [69, 98, 11, 15, 12]. This computational cost is especially prohibitive in inverse design settings, where many forward solves for optimization. Inverse design for nano-scale heat transport was first introduced in [20], where optimal material distributions were identified for various problems, such as maximum dissipation and temperature control. Subsequently, Ref. [77] developed a differentiable phonon BTE solver based on JAX [8] to design thermal metamaterials with prescribed effective thermal conductivity. This tool, which was merged in OpenBTE [80], performed density-based topology optimization [26, 62] based on the three-field approach [84] and a novel interpolation technique to map the density of the material into an effective transmission coefficient.
Despite this advance, a computational bottleneck arises in scenarios where multiple optimization runs need to create multiple target conductivities, because the topology optimization needs to be rerun from scratch for each target, which can become prohibitive due to the cost of solving the BTE.
@@
To address the computational costs, data-driven surrogates aim to accelerate the simulation and optimization of typically high-dimensional, design parameters search [25, 39]. In particular, machine learning surrogates hold the promise to learn the mapping from design parameters (i.e. material structure) to a desired low-dimensional target property (i.e. effective thermal conductivity) and dramatically reduce computational expense for the evaluation of the property. In addition, they offer a continuous relaxation of the parameters and the gradient may be computed efficiently via automatic differentiation [27, 5]. However, purely data-driven models come with the cost of the training data and are limited by the curse of dimensionality that require large training sets that grow exponentially with the number of variables [7, 30], causing an upfront, potentially large, training cost. Moreover, surrogate models may generalize poorly outside the generated data distribution, being often unreliable at extrapolation compared to interpolation [75]. Hybrid models from scientific machine learning (SciML) aim at reducing this data need by incorporating domain knowledge inside the model. Scientific and physics-informed approaches (SciML, PIML) have been used to create faster approximations [39, 72, 86, 43, 44, 42, 3, 2]. These include physics-informed networks [75, 60], solver corrections [17, 19], and operator learning [55, 47, 59]. Many of these models follow physical rules by including symmetries [18], structural patterns [61], or geometric information [10, 24], or by directly embedding solvers in the learning process  [71]. This approach has enabled major results in protein folding [38] and weather forecasting [68, 45]. Physics-Informed Neural Networks (PINNs)  [75] have been extensively applied to diffusive  [66] and ballistic heat transfer [51, 53, 52, 56] and to inverse problems in the diffusive  [13, 6, 74] and ballistic regime  [97]. Focusing on nano-scale transfer, in  [51, 53, 96, 56], PINNs are trained minimizing the sum of the residual with respect to the mode-resolved BTE and its boundary conditions at different points accross the domain (data-free ML). Recently, Ref. [97] proposed an architecture that extends to identifying some of the PDE parameters in an inverse problem setting. However, PINNs are solvers that learn a PDE solution rather than surrogate models that learn the parameterized function of a property. PINNs are typically not parameterized, but previous work introduced one physical parameter to the solver [51, 53, 96, 56] ––– the characteristic length which is hard to define and compute for complex geometries. In contrast, our surrogate model is geared towards design with twenty-five geometry parameters. Multi-fidelity (MF) DeepONets [59] learn a parameterized PDE solver from data, integrating low-fidelity simulations with sparse high-fidelity data [58]. This approach improved predictive accuracy and reduced the cost of data generation. However, it required 1000 high-fidelity data, and the low-fidelity model still required 10,000 approximate BTE solves to work, which remained computationally expensive. In contrast, we need only a few hundred high-fidelity data and use the much cheaper macroscopic approximation of Fourier equation inside our model. Note that operator learning techniques like DeepONet reduce to a single neural network (one of our baselines) when used as a parameterized surrogate model [71].
@@
We developed a SciML surrogate for the steady-state Boltzmann Transport Equation based on Physics Enhanced Deep Surrogate (PEDS) [71]. PEDS combines a low-fidelity solver, enforcing physical behavior, with a neural network that learns the low-fidelity solver input that makes it accurate for a target property. The cheap solver provides an inductive bias [67] with a relaxed version of the physics (Sec. 2). In our case, we employ a differentiable Fourier simulator as the low-fidelity approximation of the BTE. To our knowledge, this is the first successful implementation of a multifidelity approach that leverages the diffusion equation for BTE. Although when computing the thermal conductivity it has a fractional error of up to 600%, the Fourier solver is $\approx$ 2300 times faster to compute (11000 times faster considering its strong batch-parallelism) and its inductive bias accelerates training and improves the generalization in out-of-distribution regions, resulting in three to four orders of magnitude accelerations of PEDS compared to BTE. Solving one high-fidelity BTE takes around 3 minutes on 4 CPUs for our illustrative example, so training costs are dominated by data simulation costs, and our method aims at reducing the data need while retaining accuracy. Combined with active learning [70], PEDS needs only 300 data points to achieve a 5.00% target fractional error (Sec.  4), dominated by fabrication error and sufficient for PDE-constrained optimization purposes (Sec.  4.2). The proposed surrogate outperforms a purely data-driven baseline, improving the model test set fractional error of up to 76% relatively to the data-driven baseline for the same number of training points (Sec.  4). Our current inverse design pipeline enables fast PDE evaluations for thermal conductivity design tasks. We achieve an average design error of $\sim$4% on 8 example design objectives, dominated by the error of the fabrication process. Thanks to the data efficiency of PEDS and Active Learning (AL), we fully amortize the training costs with only four designs.
@@ 2 An Overview of Physics Enhanced Deep Surrogates
Figure 1: PEDS Diagram The main panel illustrates the PEDS workflow: starting from a vector encoding the parametrized topology $G$, a neural-network generator $\text{generator}_{NN}$ produces a coarser, non-linear transformation of the design space. This generated topology is then combined with a coarsified representation of the original geometry via a linear combination; the mixing coefficient $w_{\phi}\in[0,1]$ is itself learned as a function of the geometry parameterization. The resulting linear combination is passed to a computationally cheap, low-fidelity version of the physics. In our case this is the Fourier solver (depicted here with a cartoon image of Ludwig Boltzmann) which evaluates the approximate physics and outputs the target macroscopic property. The top right inset highlights how the training dataset is constructed: high-fidelity ground truth labels are produced by running expensive Boltzmann Transport Equation simulations using the OpenBTE software (represented here by a historical photograph of Ludwig Boltzmann, courtesy of the University of Frankfurt). The top left insets represent the considered baselines: a Gaussian Process with a Matern kernel and an MLP.
@@
PEDS is a scientific machine learning model comprising a computationally cheap low-fidelity solver paired with a neural network generator [71], as in Eq. 1 and illustrated in Fig. 1. The neural network applies a nonlinear transformation to the design-parameter space; the transformed parameters are passed to the low-fidelity solver, which approximates the governing physics and returns the target property. The whole architecture is trained end-to-end to predict the target property. PEDS belongs to the input-space representation machine learning field [48, 64] and is similar to neural space mapping [4] or coarse/fine grid mapping [21], with the relevant differences that the output parametrization is not the same as the initial and that in more classic input-space machine learning there is no mixing of the generated input with an input from field knowledge. The low-fidelity solver may be the same numerical method as the high-fidelity PDE solver, run at lower spatial resolution or with higher tolerance, or it may incorporate deliberate simplifications of the physics (for example, by linearizing a nonlinear term). This low-fidelity solver can produce large errors in the target output (in our case 220 % on average as it can be seen in Figure  2), but it is orders of magnitude faster than the high-fidelity model while qualitatively preserving at least some of the underlying physics (e.g. the domain boundary conditions). To incorporate further physical bias, the geometry fed to the approximate solver can be a linear combination of the generated topology and the original geometry, possibly coarsified to match the generated topology dimension. The coefficient of the convex linear combination, from now on referred to as the mixing coefficient, is learned as a function of the input geometry. The modified geometry can also enforce physically sound biases like symmetries. The neural network weights are learned contextually using backpropagation and the adjoint simulation of the PDE. PEDS has previously shown great improvements in terms of accuracy and data efficiency compared to other deep parametric models and classical surrogates in similar tasks, including surrogate modeling for diffusion and diffusion-reaction equations, and for the more complex Maxwell’s equations [71]. PEDS is defined as
@@
$\kappa\approx f_{fourier}\left(w_{\phi}\text{generator}_{NN}(G)+(1-w_{\phi})\text{downsample}(G)\right),$ (1)
@@
where $G$ parameterizes the surrogate model input geometry, $f_{fourier}$ is the low-fidelity solver, $w_{\phi}$ the mixing coefficient, $\text{generator}_{NN}$ the neural generator of the solver input, and downsample the solver input from field knowledge. PEDS fits greatly the Boltzmann Transport Equation, as the mixing coefficient $w_{\phi}$ adapts predictions between macro- and nano-scale effective conductivities, naturally capturing the smooth transition from diffusive to ballistic transport. Moreover, there is wide literature converging to a clear choice for the solver to approximate the physics: the Fourier equation, also known as heat conduction equation or Poisson equation. The Fourier Equation is the macro-scale approximation of the BTE and it can be derived through simplification of the BTE. In Fig.  2a and  2b the temperature fields of BTE and Fourier are shown for a representative geometry. Looking at the contour lines, it is easy to appreciate the difference: in the BTE solution, the temperature field exhibits clear non-diffusive effects, with strongly distorted isothermal contours around the pores that highlight the influence of boundary scattering. In contrast, the Fourier solver produces a smooth field with symmetric and evenly spaced isothermal contours. The effective thermal conductivity obtained from BTE for this case is substantially damped: 23 $W/mK$ against the Fourier solution $105W/mK$. In Fig.  2c, a comparison of the effective thermal conductivity obtained through BTE and Fourier is showcased. The Fourier field systematically overestimates the thermal conductivity when compared with the BTE results (see the line x = y for reference). The overestimation is most severe for designs where the BTE conductivity is small, corresponding to cases dominated by ballistic transport, with fractional errors exceeding 700% in the most extreme examples. When the average conductivity approaches the bulk conductivity (i.e. assuming uniformity of the material and absence of pores), the BTE scalar temperature field approaches the Fourier field. The Fourier Equation meets the inclusion criteria defined in [71] , since uniform geometries encompass the full range of effective conductivities described by the BTE.
@@
The neural network parameters and mixing coefficient are learned jointly in an end-to-end manner, making a differentiable low-fidelity solver essential. To efficiently compute the gradients of the PDE solution with respect to multiple design parameters, we exploit the adjoint simulation, also known as reverse-mode automatic differentiation [5] or backpropagation, depending on the reader’s background. Instead of using system inversion or forward numerical differentiation, gradients are obtained by solving the adjoint equation, an auxiliary PDE that captures efficiently the solution’s sensitivity to design parameters. Obtaining the gradients amounts only to solving another system similar to the forward system. For self-adjoint operators, like Fourier with periodic boundary conditions, the adjoint system is identical to the original, and further computational efficiencies can be achieved depending on the re-usability of the forward solution strategy. Training and inference were performed on Intel Core i5 quad-core processors (2.3 GHz, 2017 generation) using CPU execution only.
@@
Figure 2: Comparison between the Fourier solver and the BTE. On the left we show the qualitative difference between the temperature fields obtained with (a) the ballistic BTE and (b) the diffusive Fourier model for the same geometry and problem setup. The domain is a square of size 100 nm, containing embedded square pores of size 10 nm, with periodic boundary conditions applied along both the x and y directions, and a temperature difference imposed between top and bottom of the domain. The effective thermal conductivity $\kappa$ obtained by the Fourier temperature field is 105 W/mK, while it is 23 W/mK for the BTE field. (c) On the right we present a quantitative comparison of the computed effective thermal conductivity for 500 nanoporous designs in our test set.